OpenStudy (anonymous):

jack usually mows his lawn in 4 hrs. Marilyn can mow the same lawn in 7 hrs. How much time would it take for them to mow the lawn together? I worked it and got 3.5, but Im not sure if thats correct.

OpenStudy (anonymous):

Jack would mow half of the lawn in 2 hours, and Marilyn would mow the other half in 3.5 hours. So 2+3.5=5.5 hours.

OpenStudy (anonymous):

It should be \[\frac{7\times4}{7+4} = 2.54\]

OpenStudy (anonymous):

In 1 hr Jack can mow 1/4 of the lawn In 1 hr Marilyn can mow 1/7 of the lawn 1 1 7 + 4 11 Together in 1 hr they can mow lawn = ----- + ----- = -------- = ------- 4 7 28 28 28 6 So they would do the full lawn in ------- = 2---- or 2.55 hrs 11 11

OpenStudy (anonymous):

@waterfalling when they work together means they work together and finish the job at the same time. so your diving the lawn into two and Jack finishing first in two hours and going off and Marilyn working for 3.5 hours is not the right method

OpenStudy (anonymous):

Ah, I see, that makes sense. I had a feeling it was a little too easy haha

OpenStudy (anonymous):

no problem.............. some question trick us, don;t they☺

OpenStudy (anonymous):

They maybe only have one mower. That should be the most common case.

OpenStudy (anonymous):

if they have only one lawn mower, how can they "work together" ??? then they will hv to work separately.... since question says "How much time would it take for them to mow the lawn together?" so we have to presume that they hv two lawn mowers to be able to work together.....

OpenStudy (anonymous):

You can share the work with one lawn mower, but you will not gain time. Working together does not have to be the same as working at the same time.